OpenStudy (anonymous):

help me parameterize the curve y=x^2 into terms of t

OpenStudy (anonymous):

please

OpenStudy (amistre64):

y = r sin(t) x = r cos(t) r = sqrt(x^2 + y^2)

OpenStudy (amistre64):

but in this case x = t .... maybe

OpenStudy (anonymous):

x = t, y= t^2

OpenStudy (amistre64):

x = t y = t^2 looks to be good

OpenStudy (anonymous):

wow....i feel dumb. you guys are right, just use x=t

OpenStudy (amistre64):

doh!! i was getting around to it lol

OpenStudy (anonymous):

this is what i get for being up all night. fail. lol